OpenStudy (anonymous):

|5x+1|=2 Solve for me pleaseeeeee! :)

OpenStudy (anonymous):

5x+1 = 2 or 5x+1 = -2

OpenStudy (anonymous):

hence x= 1/5 0r x= -3/5

OpenStudy (anonymous):

aww thank you so much :)